How much water is needed to dilute a mixture of 0.25 moles NaOH and 250 ml H2O so that the solution will have a concentration of 5 ppm?


1
Expert's answer
2021-02-22T05:54:48-0500

Answer: 2 tons or 1999740g


Explanation

5ppm = 0.0005% = 5 * 10^(-6)

"5 * 10^{-6} = \frac {m_{NaOH}} {m_{NaOH} + m_{H2O} + x};"

m(NaOH) = n * M = 0.25 * 40 = 10 (g)

x = 1999740 grams, 2 tons.


SOLUTION: x = 1999740 grams, 2 tons.